OpenStudy (anonymous):

In a race, A beats B by 15 metres and C by 29 metres. When B and C run over the course together, B wins by 15 metres. The length of the course is ____

OpenStudy (anonymous):

srry 215 200 171

OpenStudy (anonymous):

200 185 171

OpenStudy (anonymous):



OpenStudy (kropot72):

Let the length of the course be x metres. When B runs x metres be gains 15 metres on C. Therefore B gains 1 metre for each x/15 metres that B runs. When B has run x - 15 metres he is 14 metres ahead of C. Therefore we can write: \[\frac{x-15}{\frac{x}{15}}=14\ .......(1)\] Solving equation (1) gives x = 225 metres.

OpenStudy (kropot72):

Therefore B gains 1 metre over C for each x/15 metres that B runs.*

OpenStudy (anonymous):

great!:-)

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

confusing slightly

OpenStudy (anonymous):

k got it!

OpenStudy (kropot72):

Good on you!